OpenStudy (anonymous):

hello! can someone help me with this assignment? s^2+3s-10=0

OpenStudy (anonymous):

answer or method?

OpenStudy (anonymous):

this one factors as \[(x+5)(x-2)=0\]

OpenStudy (anonymous):

so \[s+5=0\iff s=-5\] \[s-2=0\iff s=2\]

OpenStudy (anonymous):

either thanks